Find the transfer function of the given bode plot if K = 10, ωί = 15, Note: The horizental axis is not linear = 150, a 3 = 200, and a4 = 632. All units are in rads. H(dB) 2 3 4 H(s) =
Answer #1

The Bode magnitude plot is, H(dB) ------ -40 dB/dec - --- -I- - OT wz wz C wz wa From the bode diagram, -40 dB/dec slope reprThe magnitude at o=150 rad/s is 10 dB. K 1+1 10 = 20log 10- K,(2) 20 (101) (8) K =(595)(3.1623) K = 199.62 The transfer funct